# Digital Gatekeeping: An Audit of Search Engine Results shows tailoring of queries on the Israel-Palestine Conflict

### v1

@@ 1. Introduction
In this paper, we aim to address this gap by presenting the first empirical study on the effects of search engine customization on global controversial topics. Using the contemporary Israel-Palestine conflict as a case study and focusing on three popular search engines—DuckDuckGo, Google, and Yahoo—we explore three main research questions: 1) Do search engine results show higher levels of customization for more polarizing queries compared to non-sensitive topics, and what factors drive this tailoring? 2) How different are the type and content of URLs provided by the search engines for conflict-related searches compared to general, non-sensitive topics? 3) Do the suggested contents vary according to geographical location and do they carry particular conflict-related leanings? Our choice of this conflict was motivated not only by the gravity of this polarizing global affair but also because it involves stakeholders across different political, geographical, cultural, and linguistic boundaries. This diversity offers a particularly rich context to examine how factors like user location, language, and browser history influence search engine customization.
@@
Performing this study presents several non-trivial challenges. Search engines operate mostly as black boxes, making it difficult to systematically examine how their algorithms function. Mimicking realistic user behavior across different regions, languages, and browsing habits without introducing bias is also essential. Other challenges include the need to distinguish between neutral and politically charged queries to draw meaningful conclusions, and define reliable metrics to measure not only the variations in search engine results but also their qualitative differences. This paper addresses these challenges and makes the following key contributions:
@@
1. Crawling method with incremental profiling: We developed an automated crawling approach to perform online searches using software bots. These bots were progressively made more “human-like” by expanding their profiles: Type 1 bots varied only by location, Type 2 bots included different browser languages, and Type 3 bots incorporated an initial online browsing period to collect cookies. We deployed these bots in various countries and directed them to the search engines with identical queries. These queries included both conflict-specific topics (e.g., “Hamas”) and general, non-specific topics (e.g., “how to tie a tie”). In this way, we maintained control over each feature’s contribution to customization while avoiding the use of private or sensitive user data.
@@
2. Metrics for measuring customization: After deploying the bots, we analyzed whether the results varied and how these variations were reflected in differences in content or conflict-related bias. To do this, we adopted a metric system based on Rank-Biased Overlap (RBO) (Moffat and Zobel, 2008), to emphasize top results, and employed additional methods to assess qualitative differences. This framework enables us to study how individual features combine to influence search engine results and serves as a tool to evaluate accuracy and potential biases beyond the Israel-Palestine case study.
@@ 2. Background and Related Work
Search engine algorithms can help shape individual views on particular topics. For example, researchers have shown that even small changes in rank biased towards a political candidate can shift the voting preferences of undecided voters by 20% (Epstein and Robertson, 2015). Search engines apply various criteria to determine the information presented online, such as keyword matching, site authority, context, location, and language (Google, 2024b; Moz, 2024). Still, the factors used to tailor results to a specific user, remain uncertain, despite their potentially large impacts (Pariser, 2011). In fact, it is often impossible to distinguish if differences are due to personalization, here defined as search results curated particularly for a user in question, considering the user’s tastes and specific profile or to customization, which, may result from differences in less personal features, such as location or language.
This work focuses on how different features can impact the order, the content, and the leaning of search-engine results, especially when a user is looking for geographical and politically sensitive information, as is the case for the Israel-Palestine conflict.
@@ 2.1. Search Engine Customization
Google is the most widely used search engine, and claims that results may differ from user to user based on “time, context and personalization” (Google, 2023). Time differences arise from data being updated at various speeds across data centers. Context includes aspects such as location, language, device type (screen, mobile or desktop), and related results, which may appear if a user engages with specific content on the results page. Personalized results —- specifically tailored to an individual user profile — are tied to a Google Account.
Yahoo offers a search engine, but its primary business focus lies on email and news services, with these being particularly popular (SimilarWeb, 2024). It claims to use past browsing history and search queries across all its recommendation systems, like email and news (Yahoo, 2024). Additionally, Yahoo email users can receive personalized search content based on their activity in these cross-platform services.
@@ 2.2. Search Engine Audits
Given the proprietary nature of search engine algorithms, researchers have relied on external audits to evaluate the extent and possible negative consequences of search engine customization. In 2013, Hannak et al. (Hannak et al., 2013) introduced a methodology based on comparing the results of a “real” person with those of a “fresh” profile for the same query. They showed that “location” and “being logged-in” were the most important factors to induce search result differences. This work only audited Google but was later expanded to include Bing, which, on average, showed more personalized results (Hannák et al., 2017).
Due to its importance, most studies focused on political customization in the context of electoral processes, and almost solely audit Google (Le et al., 2019; Robertson et al., 2018b, a; Krafft et al., 2019; Diakopoulos et al., 2018). These often examine the effects of different factors in isolation (typically location or browsing history), and have produced mixed findings. While some earlier research found no significant personalization in search results based on users’ political profiles (Krafft et al., 2019; Robertson et al., 2018a; Haim et al., 2017), more recent studies suggest that a user’s political leaning can influence the political bias of results, especially in top news listings (Le et al., 2019; Robertson et al., 2018b). Moreover, location has been shown to have a stronger influence on local queries than on political ones (Kliman-Silver et al., 2015). Importantly, and unlike our study, this literature primarily compares URLs and only to a smaller extent the differences in the content that these URLs direct to. Important exceptions are the work of Le et al. (Le et al., 2019), measuring the leaning of search engine results page domains, but not of individual URLs, Robertson et al.(Robertson et al., 2018b), which groups search results in categories, and Hu et al. (Hu et al., 2019), that scores partisan bias of results’ snippets and the corresponding web pages.
@@ 3. Methodology
Our methodology relies on a Web crawling system, developed to explore how search engine results may change for different user profiles. It enables the execution of Web crawlers (bots) that can mimic typical online user behavior, be deployed to different locations, and make queries that are either related to the Israel-Palestine conflict or more general, non-sensitive queries. Figure 1 summarizes our methodology for developing and implementing this querying and auditing system. It is divided into three main stages: 1) design of each bot, including its incremental feature design and deployment, 2) audit of search engine results, and 3) evaluation of search result differences in terms of URLs, content, and possible leaning. All code and results will be made available upon acceptance of the manuscript. Next, we explain these stages.
@@ 3.2. Search Engine Audit
In our study, three search engines were audited - DuckDuckGo, Google and Yahoo. Google was chosen for being the most used search engine worldwide; DuckDuckGo for its emphasis on privacy; and Yahoo for its association with a popular email service and extensive claims of personalization.
@@ 3.2.2. Search engine queries:
The total of 27 queries per group (general and conflict-specific) were executed during the search engine audits for both Type 1 and Type 2. The audit deploying Type 3 bots performed a subset of 10 queries of the original set (Appendix, Table S1, text in bold). To maintain statistical integrity, each comparison of results ensured a consistent number of bots (accounting for occasional IP failures, Appendix S2) and queries per location (Appendix S2).
@@ 3.3.1. Quantifying Differences in URLs:
where $S$ and $T$ are two ranked lists, $d$ is the rank and $p$ determines the top-weightiness of the metric. For our analysis, $p$ was set to 0.7, meaning that the first 10 results account for 99% of the total value of the calculated RBO.
The differences between two lists of search engine URLs results were calculated as $D=1-RBO$. Other metrics, such as edit distance and intersections of lists were also applied and gave similar results.
@@ 3.3.2. Quantifying Differences in Content
To account for the possibility that the bots are receiving similar results through different local or language specific URLs (e.g.wikipedia.il vs. wikipedia.br), we classified all domains encountered (by the three types of bots, for both type of queries) in broad categories (e.g. Entertainment, News, etc.). The classification was initially done automatically by ChatGPT-4o, and then verified by one of the researchers for all domains (Appendix, Table S9). After this classification was performed, the same measurement, $D=1-RBO$ was computed, but considering lists of websites categories ([[News_1, Political_1, …] vs. [Political_1, News_1…]), instead of the URL lists.
@@ 3.3.3. Quantifying Differences in Leaning
To quantify whether the bots received results with particular leanings regarding the conflict, we focused on the URLs from domains previously classified as “News”, for the 10 searches that were common to bots of all three types (see Appendix, Table S1). These articles were scraped, all domain-related information removed, and their title and content collected and automatically translated to English, to minimize possible coder bias from knowing the language or the news source.
@@
3.4. Ethics
This study, ethically approved by the European Union (funder), deliberately used automated bots to avoid personal data use. The main risk — supporting malicious websites through page views and ad revenue during bot training — was deemed minimal because the bots mainly visited general and news sites, and the scale of our visits was very small compared to the sites’ total traffic. The scientific and social benefits of this research outweigh these concerns.
@@ 4. Results
We organized the results as answers to the main questions guiding our work. We ask if our bots (or proxy-users) see different results when querying search engines and, if so, how user’s features affect those differences, particularly for sensitive topics such as the ongoing Israel-Palestine conflict (section 4.1). This was done incrementally, by varying the level of bot customization: first varying only the location (Type 1), then varying both location and browser language (Type 2), and finally varying location, browser language, and browsing history (Type 3). Next, we tested whether disparities in result URLs also correspond to meaningful differences in content (section 4.2) and/or content leaning (section 4.3).
@@ Acknowledgements.
HoneyComb, BrightData

### v2

@@ Abstract.
* current affiliation: Portuguese National Cybersecurity Centre, Lisbon, Portugal.
@@ 1. Introduction
In this paper, we aim to address this gap by presenting the first empirical study on the effects of search engine customization on global controversial topics. We focus on the contemporary Israel-Palestine conflict and analyse the search results offered by three popular search engines—DuckDuckGo, Google, and Yahoo. We chose to study this conflict not only because of the seriousness of this polarizing global affair but also because it involves stakeholders across different political, geographical, cultural, and linguistic boundaries. This diversity offers a particularly rich context to examine how factors like user location, language, and browser history influence search engine customization. We explore three main research questions: 1) Do search engine results show higher levels of customization for more polarizing queries compared to non-sensitive topics, and, if so, what factors drive this tailoring? 2) How different in type and content are the URLs provided by the search engines for conflict-related searches when compared to general, non-sensitive topics? 3) Do the suggested contents vary according to geographical location and, if so, do they carry particular conflict-related leanings?
@@
Performing this study presented several non-trivial challenges. First, search engines operate mostly as black boxes, making it difficult to systematically examine how their algorithms function. Second, it is essential to mimick realistic user behavior across different regions, languages, and browsing habits without introducing bias. Third, it is neither trivial to distinguish between neutral and politically charged queries to draw meaningful conclusions, and define reliable metrics to measure not only the variations in search engine results. This paper addresses these challenges and makes the following key contributions:
@@
1. Crawling method with incremental profiling: We developed an automated crawling approach to perform online searches using software bots. These bots were progressively made more “human-like” by expanding their profiles: Type 1 bots varied only by location, Type 2 bots included different browser languages, and Type 3 bots incorporated an initial online browsing period to collect cookies. We deployed these bots in various countries and directed them to the search engines to perform identical queries. These queries included both conflict-specific topics (e.g., “Hamas”) and general, non-specific topics (e.g., “how to tie a tie”). In this way, we maintained control over each feature’s contribution to customization while avoiding the use of private or sensitive user data.
@@
2. Metrics for measuring customization: After deploying the bots, we analyzed whether the results varied and whether these variations were reflected in differences in content or in conflict-related bias. To do this end, we adopted a metric system based on Rank-Biased Overlap (RBO) (Moffat and Zobel, 2008), which emphasizes top results, and employed additional methods to assess qualitative differences (details below). This framework enables us to study how individual features combine to influence search engine results and serves as a tool to evaluate accuracy and potential biases beyond the Israel-Palestine case study.
@@ 2. Background and Related Work
Search engines apply various criteria to determine the information presented online, such as keyword matching, site authority, context, location, and language (Google, 2024b; Moz, 2024). These algorithms choices can help shape individual views on particular topics, and studies have shown that even small changes in rank biased towards a political candidate can shift the voting preferences of undecided voters by 20% (Epstein and Robertson, 2015). As the factors used to tailor results to a specific user, remain uncertain, despite their potentially large impacts (Pariser, 2011), it is fundamental to develop independent tools to explore these choices. In fact, it is often impossible to distinguish if differences are due to personalization, here defined as search results curated particularly for a user in question, considering the user’s tastes and specific profile or to customization, which, may result from differences in less personal features, such as location or language.
This work focuses on how different features can impact the order, the content, and the leaning of search-engine results, especially when a user is looking for geographical and politically sensitive information, as is the case of the Israel-Palestine conflict.
@@ 2.1. Search Engine Customization
Google is the most widely used search engine, and claims that results may differ from user to user based on “time, context and personalization” (Google, 2023). Time differences arise from data being updated at various speeds across data centers. Context includes aspects such as location, language, device type (screen, mobile or desktop), and related results, which may appear if a user engages with specific content on the results page. Personalized results - specifically tailored to an individual user profile — are claimed to be tied to a Google Account.
Yahoo offers a search engine, but its primary business focus lies on email and news services, with these being particularly popular (SimilarWeb, 2024). It openly claims to use past browsing history and search queries across all its recommendation systems, including email and news (Yahoo, 2024). Additionally, Yahoo email users can receive personalized search content based on their activity whithin these cross-platform services.
@@ 2.2. Search Engine Audits
Given the proprietary nature of search engine algorithms, researchers have relied on external audits to evaluate the extent and possible negative consequences of search engine customization. In 2013, Hannak et al. (Hannak et al., 2013) introduced a methodology based on comparing the results of a “real” person with those of a “fresh” profile for the same query. They showed that “location” and “being logged-in” were the most important factors to induce search result differences. This work only audited Google but was later expanded to include Bing, which, on average, showed even more personalized results (Hannák et al., 2017).
Due to its importance, most past studies focused on political customization in the context of electoral processes, and almost solely audit Google (Le et al., 2019; Robertson et al., 2018b, a; Krafft et al., 2019; Diakopoulos et al., 2018). These often examine the effects of different factors in isolation (typically either location or browsing history), and have produced mixed findings. While some earlier research found no significant personalization in search results based on users’ political profiles (Krafft et al., 2019; Robertson et al., 2018a; Haim et al., 2017), more recent studies suggest that a user’s political leaning can influence the political bias of results, especially in top news listings (Le et al., 2019; Robertson et al., 2018b). Moreover, location has been shown to have a stronger influence on local queries than on political ones (Kliman-Silver et al., 2015). Importantly, and unlike our study, this literature primarily compares URLs, and only to a smaller extent, it investigates differences in the content that these URLs direct to. Important exceptions are the work of Le et al. (Le et al., 2019), measuring the leaning of search engine results page domains, but not of individual URLs, of Robertson et al.(Robertson et al., 2018b), that groups search results in categories, and Hu et al. (Hu et al., 2019), that scores partisan bias of results’ snippets and the corresponding web pages.
@@ 3. Methodology
Our methodology relies on a Web crawling system, specifically developed to explore how search engine results may change for different user profiles. It enables the execution of Web crawlers (bots) that can mimic typical online user behavior, be deployed to different locations, and make queries that are either related to the Israel-Palestine conflict or more general, non-sensitive queries. Figure 1 summarizes our methodology for developing and implementing this querying and auditing system. It is divided into three main stages: 1) design of each bot, including its incremental feature design and deployment, 2) audit of search engine results, and 3) evaluation of search result differences in terms of URLs, content, and possible leaning.
@@ 3.2. Search Engine Audit
In our study, three search engines were audited - DuckDuckGo, Google and Yahoo. Google was chosen for being the most used search engine worldwide; DuckDuckGo for its emphasis on privacy; and Yahoo for its association with a popular email service and extensive claims of extensive personalization.
@@ 3.2.2. Search engine queries:
A total of 27 queries per group (general and conflict-specific) were executed during the search engine audits for both Type 1 and Type 2. The audit deploying Type 3 bots performed a subset of 10 queries of the original set (Appendix, Table S1, text in bold). To maintain statistical integrity, each comparison of results ensured a consistent number of bots (accounting for occasional IP failures, Appendix S2) and queries per location (Appendix S2).
@@ 3.3.1. Quantifying Differences in URLs:
where $S$ and $T$ are two ranked lists, $d$ is the rank and $p$ determines the top-weightiness of the metric. For our analysis, $p$ was set to 0.7, meaning that the first 10 results account for 99% of the total value of the calculated RBO. The differences between two lists of search engine URLs results were calculated as $D=1-RBO$. Other metrics, such as edit distance and intersections of lists were also applied and gave similar results.
@@ 3.3.2. Quantifying Differences in Content
To account for the possibility that the bots are receiving similar results through different local or language specific URLs (e.g.wikipedia.il vs. wikipedia.br), we classified all domains encountered (by the three types of bots, for both types of queries) in broad categories (e.g. Entertainment, News, etc.). The classification was initially done automatically by ChatGPT-4o, and then manually verified by one of the researchers for all domains (Appendix, Table S9). After this classification was performed, the same measurement, $D=1-RBO$ was computed, but considering lists of websites categories ([[News_1, Political_1, …] vs. [Political_1, News_1…]), instead of the URL lists.
@@ 3.3.3. Quantifying Differences in Leaning
To quantify whether the bots received results with particular leanings regarding the conflict, we focused on the URLs from domains classified as “News”, for the 10 searches that were common to bots of all three types (see Appendix, Table S1). These articles were scraped, all domain-related information removed, and their title and content collected and automatically translated to English, to minimize possible coder bias from knowing the language or the news source.
@@ 4. Results
We organized the results as answers to the main questions guiding our work. We ask if our bots (or proxy-users) see different results when querying search engines and, if so, how user’s features affect those differences (section 4.1). This was done incrementally, by varying the level of bot customization: first varying only the location (Type 1), then varying both location and browser language (Type 2), and finally varying location, browser language, and browsing history (Type 3). Next, we tested whether disparities in result URLs also correspond to meaningful differences in content (section 4.2) and/or content leaning (section 4.3) towards Israel or Palestine.
@@ Acknowledgements.
We thank members of the Social Physics and Complexity (SPAC) group at LIP for comments and critical reading of the manuscript. We thank HoneyComb for support in identifying meaningful topics for this study and BrightData for generously providing proxy services, free of charge. This research was partially funded by ERC Stg FARE (853566) and ERC PoC FARE_Audit (101100653), both to JGS, and by FCT PhD fellowship (2022.12547.BD) to ID.